and you see a child on top of a roof of a building kick a soccer ball. The soccer ball is kicked at 51° from the edge of the building with an initial velocity of 17 m/s and lands 67 meters away from the wall. How tall, in meters, is the building that the child is standing on?

A boy uses a slingshot to launch a pebble straight up into the air. The pebble reaches a height of 37.0 m above the launch point 0.8 seconds later. Assume air resistance is negligible.
Answer #1

(1)

Given: \theta = 51° , vi = 17 m/s , x = 67 m

Goal: Height of the building (h)

Solution:

Let the height of the building be h.

Then by the trajectory equation , we have :

y = h + xtan\theta - gx2/(2v2cos2\theta)

0 = h + (67)tan(51) -[(9.81)(67)2/2(17)2cos2(51)]

By solving , we get

h = 109.64 m